# Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector

### v1

@@ 5.5.2 In-Pixel Detection Efficiency
Building on previous beam test results [13], we further probed the in-pixel detection performance by using the attenuation-tunable infrared laser as an accessible proxy for detection efficiency and charge collection efficiency in response to external excitation.
@@
As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s detection efficiency in response to it to mimic the pathways of charged particles. An RS scan was performed over 100,000 frames (40 seconds) under a 25 Hz laser operation for each reverse bias value across the four in-pixel positions. Given that 1000 laser pulse triggers are expected during this period, the detection efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses. Results shown in Fig. 16 demonstrate that applying a reverse bias from 0 to –2 V had already enhanced the detection efficiency at in-pixel positions A, B, and C to nearly 100%, underscoring the performance improvement enabled by the reverse bias. These observations are in agreement with beam tests of other MAPS sensors [23], providing an imperative reference to further beam test results.
@@
Furthermore, a laser with tuning method was developed and performed in order to quantitatively characterize the sensor’s charge collection capability. The most straightforward way would be to measure the charge deposit in the sensor under external excitation, however, JadePix-3 employs a binary digital readout, and it lacks the analog output of the collected charge. Therefore, an external probe was required whose intensity can be continuously varied to infer the pixel’s charge collection sensitivity. By using the laser’s tunable attenuation as a proxy for the amount of injected charge, the sensor’s binary readout can be translated into a quantitative efficiency index in a method described as follows. The laser frequency was set to 25 Hz, and the chip was operated with 20,000 Rolling Shutter frames, equivalent to an 8-second acquisition window. Under these conditions, if every laser pulse generates enough charge that exceeds the detection threshold, 200 hits would be recorded by the pixel. The laser pulse energy was then attenuated by incrementally increasing the tune value until the number of recorded hits dropped to around 100333The actual experimental values are not exactly 100 and range between 90 and 110 due to unavoidable fluctuation.. The tune value at this point, designated as the Detection Sensitivity Index (DSI), was recorded for each of the four marked positions. This DSI is defined as the laser tune value at which the number of detected hits drops to approximately 100 out of 200 possible triggers; a higher DSI indicates that the pixel remains responsive even at lower laser intensities, thus signifying more effective charge collection efficiency and thus improved detection performance.
Figure 17: Detection Sensitivity Index (DSI) as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 15.
The variation of the DSI with reverse bias $V_{\rm sub}$ was measured and is depicted in Fig. 17. The resulting curves reveal distinct trends across the four in-pixel positions, providing more intricate information about the in-pixel variance. Position A, the pixel center, consistently exhibited the highest DSI, indicating peak detection efficiency, whereas Position D showed the lowest DSI, reflecting diminished sensitivity at the pixel boundary. Position B (long-edge center) and C (short-edge center) displayed intermediate DSI, with Position B substantially outperforming Position C. This aligns with findings from the JadePix-3 beam test measurement, which revealed notably higher detection efficiency at the long-edge center (Position B) compared to the short-edge center (Position C), where efficiency was significantly lower and closely resembled that of the pixel corner (Position D). Notably, the DSI increased with increasing reverse bias across all positions, suggesting enhanced charge collection efficiency at higher $V_{\rm sub}$ due to the expanding depletion volume.
@@ 5.6 Optimal Reverse Bias Setting
A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that a reverse bias of approximately –5 V offers the best overall trade-off. This optimal working-point not only improves detection efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
@@ 6 Conclusions and Outlook
To further characterize the performance of JadePix-3, a prototype MAPS developed for the CEPC vertex detector, detailed laboratory-based measurements were performed, in particular with substrate reverse bias voltages applied. A comprehensive test workflow was developed and implemented for the systematic characterization of JadePix-3. Experimental results show that the JadePix-3 sensor can work stably when a substrate reverse bias is applied. Moreover, the reverse bias has a considerable influence on the performance of the sensor, typically being beneficial. Our results reveal that applying reverse bias substantially enhances sensor performance, reducing input capacitance, enlarging the depletion region, and improving charge collection efficiency, while also lowering fake-hit rates. Optimal reverse bias setting is also determined within this integrated test procedure. This test methodology not only confirms JadePix-3’s potential for further performance improvement but also establishes a robust methodology for the evaluation and optimization of similar MAPS devices in future particle tracking applications.
This result acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test. Moreover, this work will serve as the architectural and procedural foundation for the tuning, and optimization of the next iteration, the JadePix-5 sensor, thereby driving forward MAPS development to meet CEPC’s requirements.

### v2

@@ 3.5.2 In-Pixel Detection Efficiency
Previous beam test results [6] indicated a non-uniform in-pixel detection efficiency, particularly between the long and short edges of the pixel. To investigate this observation, we further probed the in-pixel performance by using an attenuation-tunable infrared laser as an accessible proxy for detection efficiency and charge collection efficiency.
@@
As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s detection efficiency in response to it to mimic the pathways of charged particles. An RS scan was performed over 100,000 frames (40 seconds) under a 25 Hz laser operation for each reverse bias value across the four in-pixel positions. Given that 1000 laser pulse triggers are expected during this period, the detection efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses.
The results shown in Fig. 14 reveal distinct trends across the four in-pixel positions, providing information about the in-pixel variance. Position A (the pixel center) consistently exhibits the highest efficiency, while Position D (the corner) shows the lowest, reflecting diminished charge collection efficiency at the pixel boundary. Position B (long-edge center) and C (short-edge center) display intermediate efficiencies. Notably, the efficiency at Position B is higher than at Position C. This aligns with findings from the JadePix-3 beam test measurement [6], which also revealed a higher detection efficiency at the long-edge center compared to that at the short-edge center. Furthermore, the detection efficiency increases significantly with reverse bias across all positions. Applying a bias of just –2 V is sufficient to raise the efficiency at the center and edge positions (A, B, and C) to nearly 100%, underscoring the performance improvement enabled by the expanding depletion volume. It is also noteworthy that an efficiency drop is observed at Position D when the bias increases from -5 V to -6 V. This may be caused by a smaller lateral depletion region under such a high bias condition, highlighting the importance of carefully selecting the operational bias voltage.
@@ 3.6 Optimal Reverse Bias Setting
A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that, under the tuning method presented in this paper, a reverse bias of approximately –5 V offers the best overall trade-off. This optimal working-point not only improves detection efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
@@ 4 Conclusions and Outlook
To further characterize the performance of JadePix-3, a prototype MAPS developed for the CEPC vertex detector, detailed laboratory-based measurements were performed, in particular with substrate reverse bias voltages applied. A systematic test workflow was implemented for the characterization of JadePix-3. Experimental results show that the JadePix-3 sensor can work stably when a substrate reverse bias is applied. Moreover, the reverse bias has a considerable influence on the performance of the sensor, typically being beneficial. Our results reveal that applying reverse bias substantially enhances sensor performance, reducing input capacitance, improving charge collection efficiency, and lowering fake-hit rates, though care must be taken as performance degradation, such as a drop in efficiency, was observed at the highest tested bias of -6 V. An optimal reverse bias setting is also determined within this test procedure. This test methodology confirms JadePix-3’s potential for further performance improvement.
These results acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test. Moreover, this work will serve as the architectural and procedural foundation for the tuning and optimization of the next sensor in the JadePix series, thereby driving forward MAPS development to meet CEPC’s requirements.